The attendance at a rock concert is 6700 people. The tickets for this concert cost $40 for floor seats and $25 for all other seats. The total income of sales was 185,500. write a linear system that models this situation. Using Cramer's rule, substitution method, elimination method, solve this linear system. Then compare all three methods and choose which method you perfer.
Work I've done
40x+25y=185,500
I need help to create the other equation for this linear system

Answer by rfer(16322)

f+a=6700
a=6700-f
----------------
40f+25a=185500
-------------
40f+25(6700-f)=185500
40f-25f+167500=185500
15f=18000
f=18000/15
f=1200 floor
---------------------
a=6700-1200
a=5500 all others
---------------
40*1200+25*5500=
48000+137500=185500